Sunday, 28 April 2013

Maintenance Update

Tues 23rd (St George's Day)
With such brill sunshine, I just had to go for a bike ride!  Guess where I ended up ...
I was finishing off one of the boot scrapers (which is ordered) and decided to stick the person's name on it using a bit of blue tack.  There was a small piece on the notice board in the TPO, so I tried to pull it off.  I was wearing those blue nitrile disposal gloves ... which are quite elastic.  I caught one finger under the other as I pulled the tack off, and there was a "twang" followed by a "ching" and a "plop!".  I had managed to catapult the blob of blue tack across the TPO, hitting the lamp over the workbench, and bouncing off that straight into my cup of tea!  I drank the tea and rescued the tack, of course.

Oh, back to 2807: Bruce's fix to the T-valve on the steam fountain didn't work.  Steam leaking quite badly on Sunday.  Loco Red Carded again.

Weds 24th.
It transpired that it was not the "valve" that was at fault, but the "nipple" on the end of the copper pipe, which was badly brazed on.  Geof and Bruce had a play with this: Geof removed it; he & Bruce prep'd it, and Carpo re-brazed it.  Geof & Bruce cleaned up the pipe and Geof refitted it.

The problem with this component is that it can only be tested when the loco is in steam.  We can hardly light up 2807 just to test one joint!  However, we have rather more confidence now, and hopefully all will be well for the F&D on Friday!

Gil was preparing stuff for the stall over the (1940s) weekend, and then angle-ground some rail off-cuts as sales items.

I was doing lettering on rail chairs and fitting brushes.  As I left, Bruce had begun painting more chairs ... so I hope we sell some over the weekend!

Thurs 25th
I popped down lunchtime-ish to check, and someone had started the warming fire; so I did a bit more boot scraping.

Fri 26th
2807 used for Fire & Drive.  Stuart preparing for the weekend stall.

Sat 27th.
No working parties this weekend because GWSR has a 1940s event.  2807 was in service. The only issue that I could see was that vacuum drops randomly during some travelling.  The suggestion is a leaking retaining valve (I think).  Whatever, Bruce had removed it, cleaned it, tested it & replaced it once already.  So it can't actually be that.

Sun 28th
2807 in service again.  Loco roster is published up to 10 May.  2807 is NOT in service during that period.


Saturday, 20 April 2013

Maintenance Update

Sat 13th.
I was otherwise engaged, so today I asked Bruce what had happened - he couldn't even remember if he was there or not!  :-))
Steve said he was.  Steve, John & others(?) certainly cleaned up a pile of rail chairs for me!

Weds 17th.
2807 had been red-carded.
The issue was that the T-valve which isolates the steam flow from the "fountain" to the condenser loops for the hydrostatic lubricator is not isolating the flow to the LHS coil. This means it cannot be shut off should that line break (i.e. would allow steam into the cab straight from the boiler, with no way of stopping it) – hence the Red card. The union beside the valve is leaking too which is why the problem with the valve was noticed. Should be just a matter of removing it, stripping it, and re-seating it.
Gil, Geof and Carpo worked on this during the morning.

Bruce re-did the inspection cap on the delivery end of the LHS injector, and cleaned the vacuum retaining valve (again).

It would appear that some of the reported issues are not clearly described, and indeed some are wrongly ascribed faults.  One does have to be careful before stripping down some component purely on the comment (and the purported fault) on the issues log!

I collected wooden wedges from Eddie (Carriage & Wagon) and played making boot scrapers all day.

Sat 20th.
Bruce removed the "pepper pot" and  measured the connection from it to the vacuum reservoir.  He's still trying to track down a potential leak there.

Gil borrowed a W-valve (sits directly below the regulator in the cab, and controls the oil flow to the cylinders).  The reported issue with ours says to re-pack it.  But you can't get at it to do so without removing the linkage connecting to the actuator from the regulator.  This appears to be a press-fit on our loco.  So, Gil was trying to work out if Carpo's was (on the 42xx).  We dare not remove ours today because 2807 is in service on Sunday.

John & Steve prepared 9 rail chairs.  This was below expectations, because the needle-gun kept running out of air!  We tried to find the problem.  We tried two different needle-guns, but they wouldn't work at all!  We replaced the front hose ... no difference; we replaced the rear hose ... no difference.  I upped the pressure in the hoses ... no difference.  So, we didn't fix it, and Steve had to keep stopping for the air to catch up with him!

Meanwhile, I painted 20 brushes and painted 10 tops plus 6 bottoms.

Anyway, the highlight of the day (for Steve & myself) was a ride in the class 45.  This "Peak" finally ran under its own power - in both directions!  We had a cab-ride out of the yard and up & down the line.  The team has spent 17 years acquiring the diesel, and then getting it going.  I had seen it many years ago in Camden sidings, after it had failed on leaving Euston.  It was one of the last Class 45s in service.  When they were first introduced (1959), the "Peaks" ran past my school, hence they are a bit nostalgic for me.  So, for the Cotswold Mainline Diesel Group it was just like us up at Llangollen, when 2807 moved under her own power.  Having had the privilege of a ride in the cab on her first trip, I shall have to buy another share!

is in service Sunday 21st, then Friday 26th (Fire & Drive) and over the 1940s weekend - Sat/Sun 27/28 April.
There is no loco roster beyond that date.


Sunday, 7 April 2013

Maintenance Update

Weds 3rd:
John Tyler was first on the scene (as usual) and he responded to my plea re cleaning rail chairs.  Then Gil arrived and diverted John onto fixing things on the loco.  Bruce also joined in, nibbling away at the list of issues raised by crews recently.

John fixed the RHS valve spindle packing as well as cleaning 8 chairs.
Dixie would have been pulled into fixing things, too, but luckily I arrived with the new "oversink handwash unit", or water heater!  This was only delivered to me at 11.00 (hence I was a tad late).  So, he spent the rest of his time wiring the water heater up. 

Bruce broke from fixing things (i.e repacking LHS injector delivery clack cap (see attached), and tightening bolts on the running board) to nip home and make a T-section of 15mm copper water pipe, which we inserted into the cold supply to feed the heater.

Gil spent much time under the loco figuring out how to fix the bottom of the ashpan (which has warped yet again).  I believe John Cruxon was bending ears about the difficulty in raking ash out of our ashpan!

Gil also measured the axlebox clearances with his long feeler gauge.

John (Daily) Mayell popped in and cleaned about 3 rail chairs and collected his new share certificate.

I finished off a bright red bootscraper (special order for a retiring fireman (bee-bah-bee-bah fireman)), repainted my Mary's birthday bootscraper (for 12th April), and had a good grind on some bolts on the rail chair stock.

Thurs 4th:
I've gone away!


Saturday 6th (info supplied by Bruce):

A bit of a mixed day.

The front part of the ashpan had distorted upwards preventing the rake from entering, and thus making it difficult to get the ash out.

On Wednesday Gilbert straightened the bottom of the ashpan, using clamps, and today Paul Stratford, assisted by Gilbert, welded it in position.

Paul also suggested that we make a spacer to hold the damper door fully open, to assist ash disposal.

Roger Deane got the water heater working and changed some of the wiring in the kitchen area.

Bruce started to replace the gland packing on the cylinder lubrication Combining Valve but hit a snag when the gland nut fouled the operating lever before unscrewing far enough to insert the new packing.

After studying the drawing and consulting Andrew Meridith it became apparent that the operating lever on the front would need to be removed, but it was stuck fast and appeared to be pressed on.

Our new extractor should pull it off, but was too big to fit into the space available, so it was abandoned until next time as it was now getting well into the afternoon.

Fred had arrived for lunch and when finished, Fred, Gilbert and Roger D decamped to Winchcombe to work on the Siphon.